# Emulating Self-attention with Convolution for Efficient Image Super-Resolution

### v1

@@ 4 Experiments
This section describes implementation details for flash attention and our extensive experiments on various tasks. For implementation details for proposed networks, please refer to Section 6 in the supplementary.
4.1 Implementation Details for Flash Attention
We aim to address the excessive memory overhead of self-attention by leveraging flash attention [7]. However, the existing flash attention implementation does not support relative positional bias [39], which is essential for performance improvements. Accordingly, we implement Flash attention leveraging flex attention [10] to incorporate relative positional bias and accelerate self-attention during inference. As shown in Table 1, our implementation significantly reduces latency and memory usage of self-attention by 16$\times$ and 12.2$\times$, respectively. As the patch size for training is 64$\times$64, we do not consider the window size above 32 as it may cause training instability due to the necessity of padding.
@@ 4.2 Classic SR Results
To evaluate our proposed networks on classic SR task, we leverage five commonly used datasets (Set5 [2], Set14 [52], B100 [33], Urban100 [20], Manga109 [34]). We measure peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) on the Y channel after cropping boundary corresponding to the upscaling factor and converting them into YCbCr space. All statistics are measured by reconstructing an HD image (1280$\times$720) image using RTX4090 GPU at FP32 precision 100 times. Memory usage is calculated by torch.cuda.max_memory_allocated function and FLOPs are measured by fvcore library. For the classic SR task, we present three variants: ESC, ESC-light, and ESC-FP. ESC and ESC-light prioritize to reduce latency, while ESC-FP focuses on reducing FLOPs and parameter size.
@@ 4.2.1 Quantitative Results on the DIV2K dataset
We compare our methods with various representative lightweight SR methods, including CNN-based models [58, 57], transformers [28, 55, 45, 62, 54, 56, 37], and state-space models [16, 17] trained on DIV2K dataset [1]. Notably, as shown in Table 2, our ESC achieves notable 0.1dB PSNR improvement on the Urban100$\times$4 dataset, while being 8.7$\times$ faster than ATD-light. Moreover, ESC-light is 1.4$\times$ faster than ELAN-light, exhibiting an impressive performance improvement of 0.34dB PSNR in the Manga109×$\times 4$ dataset. These results exhibit that state-of-the-art performance can be achieved even when most self-attentions are replaced with convolutions, suggesting replacing self-attention with convolutions does not compromise representational power. Furthermore, ESC-FP outperforms MambaIRV2-light over 0.18dB on the Urban100×$\times 4$ dataset while reducing FLOPs and parameters by 20% and 31%, respectively. This result underscores that our design is effective even when reducing FLOPs and parameter size.
@@ 4.2.2 Visual Results on the DIV2K dataset
4.2.3 Ablation Study
We conduct an ablation study to justify the proposed network’s composition, as shown in Table 3. First, we introduce two variants, one that uses only the self-attention module and another that uses only the ConvAttn module. Both variants exhibit a performance drop compared to ESC, suggesting that leveraging both self-attention and ConvAttn to balance performance and efficiency is crucial. Next, we vary the $LK$’s kernel size to 9$\times$9 and 17$\times$17 to determine the optimal size. In both cases, we observe a drop in performance, confirming that a kernel size of 13$\times$13 is optimal. Then, we evaluate performance changes by altering the convolution type used in the ConvAttn module. We observe a drop in performance in variants that neither share $LK$ nor employ input-dependent $DK$. This indicates that emulating the input-dependent weighting of self-attention is essential. Accordingly, we add $DK$ without sharing $LK$ and observe some performance improvement, though it still falls short of ESC. These results suggest that employing highly sophisticated modeling may not be necessary when using convolution to emulate self-attention’s ability to capture long-range dependencies. Instead, sharing $LK$, which reduces the number of parameters and thereby eases optimization [8, 30], appears to have a more significant impact on performance. Finally, we investigate the effect of the self-attention window size by setting it to 16 and 24 and adding extra layers. In every case, we observe a performance drop, which confirms that enlarging the window size to enhance self-attention is effective even when only a single self-attention is leveraged in each block.
@@ 4.2.4 Analysis on ConvAttn Module
To verify the proposed ConvAttn module extracts similar features as self-attention, we examine the features extracted by ConvAttn modules. Self-attentions are known to extract shape-biased structural features [38, 8], unlike 3$\times$3 convolution, which mainly extracts local texture and edges. Indeed, as shown in Figure 6, the visualized average feature maps indicate that $F^{idt}$, which passed through only the ConvFFN module with a kernel size of 3$\times$3, primarily captures textures such as hair. In contrast, the feature maps extracted using both $LK$ and $DK$ exhibit a bias towards depth and structure, indicating that the proposed ConvAttn module extracts features similar to self-attention. Additionally, we observe that the inter-layer similarity is reduced when both $LK$ and $DK$ are used, compared to using $LK$ alone This observation indicates that $DK$s are extracting diverse layer-specific features, providing robust evidence of the effectiveness of dynamic weights.
Figure 6: Visualized average feature map and inter-layer CKA similarity on the features extracted by ConvAttn. We average all ConvAttn’s outputs while restoring the Lenna in the Set14 dataset. CKA similarities are measured on the Urban100$\times$4 dataset.
@@ 4.2.5 Local Attribution Map
We use the Local Attribution Map (LAM) [15] to verify our networks can extract long-range dependencies despite using less self-attention. We compare our network with studies that aim to expand the receptive field of self-attention by using various self-attention implementations, including global-level attention [45], permuted self-attention [62], and spatial-channel correlation (which includes both window and channel-wise self-attention) [56]. Notably, as described in Figure 7, our ESC references the widest range to restore the red bounding box along with the highest diffusion index. This result exhibits our network effectively captures long-range dependencies even though most self-attentions are replaced by the ConvAttn module, suggesting that our ConvAttn module contributes to retaining the large receptive field of transformers.
4.2.6 Quantitaive Results on the DFLIP dataset
Transformers exhibit significant performance improvements when trained on large-scale datasets [11, 3]. To verify our model also retains this benefit, we train and evaluate various representative transformers – including SRFormer-light [62], ATD-light [54], and HiT-SRF [56] – alongside our ESC on a large dataset. For this purpose, we leverage DIV2K+Flicker2K+LSDIR+DiverSeg-IP (DFLIP) [1, 43, 26, 35] dataset. As illustrated in Table 4, our approach outperforms the other methods by a substantial margin across all scales and evaluation datasets. These results indicate that replacing self-attention with ConvAttn does not sacrifice the data scaling capability of transformers.
@@ 4.4 Real-World Super-Resolution
We further test our network on real-world SR tasks that aim to restore images with unknown degradations. To this end, we introduce a variant called ESC-Real and compare them to representative real-world SR methods, including RealESRGAN [47], SwinIR-Real [28], and DASR [48]. For our evaluation, we use the commonly employed RealSRSet dataset [53]. As shown in Table 9, our ESC-Real recovers more fine-grained details than the other methods while significantly reducing unpleasant artifacts in both graphic and natural images. This result demonstrates that our proposed methods are effective even when dealing with complex unknown degradations, underscoring our methods’ outstanding representational capability.
@@ 5 Conclusion
In this paper, we propose ESC, a novel lightweight SR network designed to mitigate the excessive memory overhead of transformers. Our extensive experiments demonstrate that most self-attention layers can be effectively replaced with carefully designed convolution to improve efficiency. Moreover, we introduce flash attention into lightweight SR tasks, significantly enhancing performance without complicated modifications to self-attention. As a result, we fully exploit the advantages of transformers by prioritizing improvements in their efficiency.
@@
The supplementary includes implementation details of our proposed networks and additional results on the arbitrary-scale super-resolution (SR) tasks.
6 Implementation Details
This section describes the details of the implementation of our methods. We begin by outlining the basic model configuration for each task, which includes the hyper-parameters $C$, $N$, and $M$. Here, $C$ represents the number of channels, $N$ denotes the number of $\mathrm{ESCBlock}$s, and $M$ indicates the number of $\mathrm{ConvAttn}$s. Next, we discuss additional components, such as the compositions of $U$ and $S$, among others. Finally, we present the training details, which cover the training datasets, the number of training iterations, the learning rate, optimizer configurations, the loss function used, and the training batch and patch size.
@@ 6.1 Classic SR
Three variants are introduced in the classic SR task: ESC-FP, ESC-light, and ESC. ESC-FP is a variant needed when it is necessary to reduce FLOPs and parameter size, while ESC-light and ESC are variants needed when it is necessary to reduce latency. The basic configurations for each variant are as follows: for ESC-FP, $C$, $N$, and $M$ are 48, 5, and 5; for ESC-light, they are 64, 3, and 5; and for ESC, they are 64, 5, and 5. All variants use Sub-Pixel Convolution (SPConv) [40] as $U$. ESC-light and ESC utilize the repeat function [12] as $S$ and add $F^{s}$ before the pixel shuffle of SPConv, while ESC-FP employs bicubic interpolation as $S$ and adds $F^{s}$ to the output of SPConv. Furthermore, ESC-FP employs decomposed $LK$ to further reduce FLOPs and parameter size and utilizes extra layer normalization in front of the $\mathrm{ConvFFN}$s. For training, we use the DIV2K [1] dataset, and for the data scaling experiments, we employ the DIV2K+Flicker2K+LSDIR+Diverseg-IP (DFLIP) [43, 26, 35] dataset. Training our networks lasts for 500K iterations, and the optimizer used is AdamW [32] with $\beta$s of 0.9 and 0.9 and a learning rate of 5e-4. We use L1 loss and 64 patches of size 64$\times$64 as input to train. The networks of scale ×$\times 3$ and $\times$4 are fine-tuned from the results of the $\times$2 scale. To train other methods [62, 56, 54] for data scaling experiments, we follow the training details described in their paper.
@@ 6.2 Arbitrary-scale SR
6.3 Real-world SR
For the real-world SR task, we introduce ESC-Real with a basic configuration of 64, 10, and 5, which denote $C$, $N$, and $M$, respectively. ESC-Real employs the same upsampler as RealESRGAN [47] and SwinIR-Real [28], utilizing it as $U$, and incorporates three layers of convolutions (c128k1g1c128k7g128c64k1g1) as $S$. Here, c, k and g denote channel, kernel, and group size for convolution, respectively. In this approach, $F^{s}$ is added into $F$. We use the RealESRGAN degradation model and DF2KOST [46] dataset to generate low-quality images. ESC-real is first trained for 400K iterations using L1 loss, then trained for 400K iterations with L1 loss, adversarial loss, and perceptual loss, using weight factors of 1, 0.1, and 1, respectively. The network architectures used for calculating adversarial loss and perceptual loss are the same as those used in RealESRGAN. In both phases, 48 patches of size 64$\times$64 are used as input for training.
7 Additional Results on Arbitrary-scale SR
This section exhibits additional results on the arbitrary-scale SR task. Additional quantitative results are measures on the four commonly used evaluation datasets, including Set5 [2], Set14 [52], B100 [33], and Urban100 [20]. Following previous research [25], we measure Peak Signal-to-Noise Ratio (PSNR) on the Y channel after cropping images’ boundary equivalent to the upscaling factor and converting them to YCbCr color space. Our ESC+LTE outperforms other methods on Urban100 at both seen and unseen scales, as shown in Table 6.

### v2

@@ 4 Experiments
This section demonstrates the effectiveness of our flash attention and our extensive experiments on various tasks. For implementation details for proposed networks, please refer to Section 7 in the supplementary.
4.1 Effectiveness of Flash Attention
We address the excessive memory overhead of self-attention by leveraging Flash Attention [10, 14]. Implementation details for Flash Attention are described in Section 6 in the supplementary. As shown in Table 1, our self-attention implementation accelerated by Flash Attention significantly reduces the latency and memory usage of self-attention at the window size of 32 by 16$\times$ and 12.2$\times$, respectively. We do not consider window sizes above 32, as it may cause training instability due to the necessity of padding, since the training patch size is 64$\times$64.
@@ 4.2 Classic SR Results
To evaluate our proposed networks on the classic SR task, we leverage five commonly used datasets (Set5 [3], Set14 [70], B100 [42], Urban100 [26], Manga109 [43]). We measure peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) on the Y channel after cropping the boundary corresponding to the upscaling factor and converting them into YCbCr space. All statistics are measured by reconstructing an HD image (1280$\times$720) image using an RTX4090 GPU at FP32 precision 100 times. Memory usage is calculated by torch.cuda.max_memory_allocated function and FLOPs are measured by fvcore library. For the classic SR task, we present three variants: ESC, ESC-light, and ESC-FP. ESC and ESC-light prioritize reducing latency, while ESC-FP focuses on reducing FLOPs and parameter size.
@@ 4.2.1 Quantitative Results on the DIV2K dataset
We compare our methods with various representative lightweight SR methods, including CNN-based models [78, 77], Transformers [36, 75, 56, 82, 73, 76, 47], and SSMs [20, 21] trained on the DIV2K dataset [1]. Notably, as shown in Table 2, our ESC achieves a notable 0.1dB PSNR improvement on the Urban100$\times$4 dataset, while being 8.7$\times$ faster than ATD-light. Moreover, ESC-light is 1.4$\times$ faster than ELAN-light, exhibiting an impressive performance improvement of 0.34dB PSNR in the Manga109×$\times 4$ dataset. These results exhibit that state-of-the-art performance can be achieved even when most self-attentions are replaced with convolutions, suggesting that replacing self-attention with convolutions does not compromise representational power. Furthermore, ESC-FP outperforms MambaIRV2-light over 0.18dB on the Urban100×$\times 4$ dataset while reducing FLOPs and parameters by 20% and 31%, respectively. This result underscores that our design is effective even when reducing FLOPs and parameter size.
@@ 4.2.2 Visual Results on the DIV2K dataset
4.2.3 Ablation Study
We conduct an ablation study to justify our network’s composition, as shown in Table 3. First, we introduce two variants, one that uses only the self-attention module and another that uses only the $\mathrm{ConvAttn}$. The number of blocks and layers is adjusted to match the latency on par with the ESC. Both variants exhibit a performance drop compared to ESC, suggesting that leveraging both self-attention and $\mathrm{ConvAttn}$ offers the best performance-efficiency trade-off. Next, we vary the $LK$’s kernel size to 9$\times$9 and 17$\times$17 to determine the optimal size. In both cases, we observe a drop in performance, confirming that a kernel size of 13$\times$13 is optimal. Then, we evaluate performance changes by altering the convolution type used in the $\mathrm{ConvAttn}$. First, we observe performance drops when neither $LK$ sharing nor $DK$ is used. Then, we add $DK$ without sharing $LK$ and observe some performance improvement, though it still falls short of ESC. This result suggests that addressing the optimization difficulty that comes from the massive parameter size is more critical than the layer-specific long-range modeling. Finally, we leverage only $LK$ sharing without $DK$ and observe significant performance drops, affirming that harmonizing both $LK$ and $DK$ is crucial. For the next step, we investigate the effect of the self-attention window size by setting it to 16 and 24 and adding extra layers. In every case, we observe a performance drop, which confirms that enlarging the window size to enhance self-attention is effective even when only a single self-attention is leveraged in each block.
@@ 4.2.4 Analysis on ConvAttn Module
This section verifies that the proposed $\mathrm{ConvAttn}$ operates similarly to self-attention. Self-attention is known to extract shape-biased structural features [48, 12], unlike 3$\times$3 convolution, which mainly extracts local texture and edges. Indeed, as shown in Figure 6 (a), $F^{idt}$ focuses on local textures and edges while $F^{att}$ concentrates on structural components. Next, to verify each convolution’s receptive field, we visualize positive gradients $\bar{G}^{+}=\tfrac{1}{NMC}\sum_{p=1}^{NMC}\max\bigl{(}0,\tfrac{\partial I^{SR}_{i,j}}{\partial F_{p}}\bigr{)}$ propagated from the output image’s center pixel $(i,j)$ to the inputs feature ($F$) of both $DK$ and $LK$ across $N$ blocks, $M$ modules, and $C$ channels, and corresponding diffusion index (DI). Figure 6 (b) clearly demonstrates that $LK$ extracts more long-range features than $DK$. Lastly, we visualize input-wise similarity ($S$) between averaged pool features extracted by $LK$ and $DK$. As demonstrated in Figure 6 (c), $S_{DK}$ exhibits higher input-dependent diversity than $S_{LK}$, confirming that $DK$ extracts more input-dependent features than $LK$.
Figure 6: Visualized feature map, gradients, and input-wise similarity. $F$ is averaged across entire $\mathrm{ConvAttn}$ while processing the Lenna image in Set14. $\bar{G}^{+}$ and $S$ are calculated on the Urban100, and 10 samples in front of the Urban100, respectively.
@@ 4.2.5 Local Attribution Map
We use the Local Attribution Map (LAM) [19] to verify that our networks can extract long-range dependencies despite using less self-attention. We compare our ESC with studies of various self-attention implementations that aim to reduce FLOPs while expanding the receptive field, such as global-level attention [56], permuted self-attention [82], and spatial-channel correlation [76]. Notably, as described in Figure 7, our ESC references the widest range of pixels to restore the red bounding box along with the highest diffusion index. This result exhibits that our network effectively captures long-range dependencies even though most self-attentions are replaced by the ConvAttn module, suggesting that our ConvAttn module contributes to retaining the large receptive field of Transformers.
4.2.6 Quantitative Results on the DFLIP dataset
Transformers exhibit significant performance improvements when trained on large-scale datasets [15, 5]. To verify our model also retains this benefit, we train and evaluate various representative Transformers – including SRFormer-light [82], ATD-light [73], and HiT-SRF [76] – alongside our ESC on a large dataset. For this purpose, we leverage DIV2K+Flickr2K+LSDIR+DiverSeg-IP (DFLIP) [1, 54, 34, 45] dataset. As illustrated in Table 4, our approach outperforms the other methods by a substantial margin across all scales and evaluation datasets. These results indicate that replacing self-attention with $\mathrm{ConvAttn}$ does not sacrifice the data scaling capability of Transformers.
@@ 4.4 Real-World Super-Resolution
We further test our network on real-world SR tasks that aim to restore images with unknown degradations. To this end, we introduce a variant called ESC-Real and compare them to representative real-world SR methods, including RealESRGAN+ [59], SwinIR-Real [36], and DASR [61]. For our evaluation, we use the commonly employed RealSRSet dataset [72]. As shown in Table 9, our ESC-Real recovers more fine-grained details than the other methods while significantly reducing unpleasant artifacts in both graphic and natural images. This result demonstrates that our proposed methods are effective even when dealing with complex unknown degradations, underscoring our methods’ outstanding representational capability. Quantitative results on real-world SR tasks are presented in Section 9 in the supplementary.
@@ 5 Conclusion
In this paper, we propose ESC, a novel lightweight SR network designed to mitigate the excessive memory overhead of Transformers. Our extensive experiments demonstrate that most self-attention layers can be effectively replaced with carefully designed convolutions to improve efficiency. Moreover, we introduce flash attention into lightweight SR tasks, significantly enhancing performance without complicated modifications to self-attention. As a result, we fully exploit the advantages of Transformers by prioritizing improvements in their efficiency.
Acknowledgments This work was supported by the National Research Foundation (NRF) grant funded by the Korea government (MSIT) [RS-2025-00562400].
@@
The supplementary includes implementation details of Flash Attention (FA) and proposed networks, additional results on the arbitrary-scale super-resolution (SR) tasks, quantitative results on real-world SR tasks, efficiency comparison beyond GPUs, and classic SR results on larger model sizes.
6 Flash Attention Implementation Details
FA [10] implements self-attention by fusing kernels and avoiding materializing a full score matrix ($S=QK^{T}$), thereby significantly reducing both memory footprint and latency. Although FA has been widely adopted across various domains, such as classification and generation, its application to SR tasks has been limited. In order to alleviate the self‐attention’s memory bottleneck, we attempted to integrate FA directly into SR architectures. However, we observed that naively applying FA leads to highly unstable training. We verified that this instability stems from FA’s fused‐kernel design, which blocks the use of relative positional bias (RelPos) [49, 40]. As shown in Table 6, training loss diverges when FA is employed without RelPos. A common alternative is to use Rotary Positional Encoding (RoPE) [51, 23], which rotates query and key tensors for positional conditioning and therefore remains compatible with fused kernels [66, 65]. Nonetheless, our experiments demonstrate that even this approach fails to deliver acceptable performance in the SR setting. To address these challenges, we leverage Flex Attention [14], which supports both user‐defined score modifications and FA. Our implementation not only resolves the memory bottleneck but also achieves superior performance by leveraging the 32$\times$32 window size for self-attention.
Table 6: Comparisons on performance by FA implementation.
Type FA [10] FA w/ RoPE [23] FA w/ RelPos (Ours)
PSNR/SSRM (U100$\times$2) NaN / NaN 32.76 / 0.9343 33.46 / 0.9395
7 Network Implementation Details
This section describes the details of the implementation of our methods. We begin by outlining the basic model configuration for each task, which includes the hyperparameters $C$, $N$, and $M$. Here, $C$ represents the number of channels, $N$ denotes the number of $\mathrm{ESCBlock}$s, and $M$ indicates the number of $\mathrm{ConvAttn}$s. Next, we describe additional components, such as the compositions of $U$ and $S$, among others. Finally, we present the training details, which cover the training datasets, the number of training iterations, the learning rate, optimizer configurations, the loss function used, and the training batch and patch size.
@@ 7.1 Classic SR
Three variants are introduced in the classic SR task: ESC-FP, ESC-light, and ESC. ESC-FP is a variant needed when it is necessary to reduce FLOPs and parameter size, while ESC-light and ESC are variants needed when it is necessary to reduce latency. The basic configurations for each variant are as follows: for ESC-FP, $C$, $N$, and $M$ are 48, 5, and 5; for ESC-light, they are 64, 3, and 5; and for ESC, they are 64, 5, and 5. All variants use Sub-Pixel Convolution (SPConv) [50] as $U$. ESC-light and ESC utilize the repeat function [16] as $S$ and add $F^{s}$ before the pixel shuffle of SPConv, while ESC-FP employs bicubic interpolation as $S$ and adds $F^{s}$ to the output of SPConv. Furthermore, ESC-FP employs decomposed $LK$ to further reduce FLOPs and parameter size and utilizes extra $LN$s in front of the $\mathrm{ConvFFN}$s. Lastly, hidden dimension ($h$) for kernel estimators is set to 8 for ESC and ESC-FP, while ESC-FP uses 4. For training, we use the DIV2K [1] dataset, and for the data scaling experiments, we employ the DIV2K+Flickr2K+LSDIR+Diverseg-IP (DFLIP) [54, 34, 45] dataset. Training our networks lasts for 500K iterations, and the optimizer used is AdamW [41] with $\beta$s of 0.9 and 0.9 and a learning rate of 5e-4. We use L1 loss and 64 patches of size 64$\times$64 as input to train. The networks of scale ×$\times 3$ and $\times$4 are fine-tuned from the results of the $\times$2 scale. To train other methods [82, 76, 73] for data scaling experiments, we follow the training details described in their paper.
@@ 7.2 Arbitrary-scale SR
7.3 Real-world SR
For the real-world SR task, we introduce ESC-Real with a basic configuration of 64, 10, and 5, which denote $C$, $N$, and $M$, respectively. ESC-Real employs the same upsampler as RealESRGAN [59] and SwinIR-Real [36], utilizing it as $U$, and incorporates four layers of shallow network (c128k1g1c128k7g128LeakyReLU($\alpha=0.2$)c64k1g1) as $S$. Here, c, k, and g denote channel, kernel, and group size for convolution, respectively. In this approach, $F^{s}$ is added into $F$. We use the RealESRGAN degradation model and DF2KOST [58] dataset to generate low-quality images. ESC-real is first trained for 1M iterations using L1 loss, then trained for 400K iterations with L1 loss, adversarial loss, and perceptual loss, using weight factors of 1, 0.1, and 1, respectively. The network architectures used for calculating adversarial loss and perceptual loss are the same as those used in RealESRGAN. In both phases, 48 patches of size 64$\times$64 are used as input for training.
8 Additional Results on Arbitrary-scale SR
This section exhibits additional results on the arbitrary-scale SR task. Additional quantitative results are measures on the four commonly used evaluation datasets, including Set5 [3], Set14 [70], B100 [42], and Urban100 [26]. Following previous research [32], we measure Peak Signal-to-Noise Ratio (PSNR) on the Y channel after cropping the image’s boundary equivalent to the upscaling factor and converting it to YCbCr color space. Our ESC+LTE outperforms other methods on Urban100 at both seen and unseen scales, as shown in Table 7,
Table 8: Quantitative comparisons on real-world SR.
Dataset Metrics RealESRGAN+ SwinIR-Real DASR ESC-Real
RealLQ250 [2] NIQE $\downarrow$ 4.1328 4.1779 4.7858 4.0556
MANIQA $\uparrow$ 0.3564 0.3400 0.2789 0.3553
MUSIQ $\uparrow$ 62.51 60.48 53.02 62.98
CLIPIQA $\uparrow$ 0.5437 0.5348 0.4631 0.5796
RealSRSet [72] NIQE $\downarrow$ 5.3430 5.1037 4.5931 5.0181
MANIQA $\uparrow$ 0.3988 0.3872 0.3277 0.3952
MUSIQ $\uparrow$ 64.25 63.68 58.82 64.58
CLIPIQA $\uparrow$ 0.5942 0.5921 0.5278 0.6156
RealSR [4] PSNR $\uparrow$ 24.53 24.71 25.86 24.52
SSIM $\uparrow$ 0.7484 0.7547 0.7617 0.7503
LPIPS $\downarrow$ 0.2729 0.2594 0.3113 0.2622
DISTS $\downarrow$ 0.1685 0.1609 0.1838 0.1671
FID $\downarrow$ 67.01 64.19 63.62 66.81
NIQE $\downarrow$ 4.6801 4.6465 5.9682 4.4848
MANIQA $\uparrow$ 0.3675 0.3504 0.2663 0.3799
MUSIQ $\uparrow$ 59.69 59.64 45.82 61.40
CLIPIQA $\uparrow$ 0.4903 0.4736 0.3629 0.5338
DRealSR [60] PSNR $\uparrow$ 26.59 26.52 28.40 26.76
SSIM $\uparrow$ 0.7988 0.7923 0.8302 0.79534
LPIPS $\downarrow$ 0.2818 0.2838 0.2962 0.2795
DISTS $\downarrow$ 0.1464 0.1461 0.1689 0.1503
FID $\downarrow$ 23.19 24.63 17.89 21.35
NIQE $\downarrow$ 4.7164 4.5683 6.3473 4.7006
MANIQA $\uparrow$ 0.3431 0.3275 0.2733 0.3442
MUSIQ $\uparrow$ 35.27 34.62 28.63 35.21
CLIPIQA $\uparrow$ 0.5179 0.5039 0.3843 0.5564
9 Quantitative Results on Real-world SR
Taking a step further, we report quantitative results for the real-world SR task. To this end, we measure a variety of metrics (PSNR@Y, SSIM, LPIPS [74], DISTS [11], FID [24], NIQE [44], MANIQA [64], MUSIQ [27], and CLIP-IQA [57]) on multiple datasets (RealLQ250 [2], RealSRSet [72], RealSR [4], and DRealSR [60]). As shown in Table 8, ESC-Real achieves the highest CLIPIQA scores on all datasets, demonstrating its ability to produce reconstructions with superior perceptual quality.
Table 9: Comparisons of latency on MacBook M2 air, and iPhone 12
Methods ELAN-lt OmniSR ASID-D8 HiT-SRF ESC-lt ESC
M2Air ($X\in\mathbb{R}^{128\times 128\times 3}$) 318.51 Failed Failed 88145.85 124.07 181.16
iPhone12 ($X\in\mathbb{R}^{32\times 32\times 3}$) 38.12 Failed Failed OOM 25.57 42.78
10 Efficiency Comparisons beyond GPUs
In real‐world deployment scenarios, networks often run on devices with far more constrained resources than GPUs. To evaluate our method under such conditions, we benchmarked several Transformer‐based SR models (ELAN [75], OmniSR [56], ASID‐D8 [47], HiT‐SRF [76]) against our ESC(‐lt) on a MacBook Air M2 and an iPhone 12. As detailed in Table 9, whereas the other Transformers either fail to compile or incur out-of-memory errors, ESC‐lt achieves up to a 61% reduction in latency compared to ELAN-light, demonstrating its efficiency in real‐world deployments.
Table 10: Comparisons of larger classic SR methods (Params$>$10M). PT denotes pre-training with 64$\times$64 patches and FT denotes fine-tuning with 96$\times$96 patches.
Method Scale #Params (M) PSNR / SSIM
Set5 Set14 B100 Urban100 Manga109
SwinIR [36] ×$\times 2$ 11.8 38.42/0.9623 34.46/0.9250 32.53/0.9041 33.81/0.9433 39.92/0.9797
EDT-B [33] 11.5 38.63/0.9632 34.80/0.9273 32.62/0.9052 34.27/0.9456 40.37/0.9811
CAT-A [9] 16.5 38.51/0.9626 34.78/0.9265 32.59/0.9047 34.26/0.9440 40.10/0.9805
ART [71] 16.4 38.56/0.9629 34.59/0.9267 32.58/0.9048 34.30/0.9452 40.24/0.9808
ACT [67] 46.0 38.46/0.9626 34.60/0.9256 32.56/0.9048 34.07/0.9443 39.95/0.9804
SRFormer [82] 10.5 38.51/0.9627 34.44/0.9253 32.57/0.9046 34.09/0.9449 40.07/0.9802
ESC (PT) 12.5 38.52/0.9626 34.57/0.9257 32.58/0.9045 34.24/0.9450 40.18/0.9803
ESC (FT) 12.5 38.59/0.9630 34.70/0.9259 32.61/0.9052 34.49/0.9466 40.38/0.9809
SwinIR [36] ×$\times 3$ 11.9 34.97/0.9318 30.93/0.8534 29.46/0.8145 29.75/0.8826 35.12/0.9537
EDT-B [33] 11.7 35.13/0.9328 31.09/0.8553 29.53/0.8165 30.07/0.8863 35.47/0.9550
CAT-A [9] 16.6 35.06/0.9326 31.04/0.8538 29.52/0.8160 30.12/0.8862 35.38/0.9546
ART [71] 16.6 35.07/0.9325 31.02/0.8541 29.51/0.8159 30.10/0.8871 35.39/0.9548
ACT [67] 46.0 35.03/0.9321 31.08/0.8541 29.51/0.8164 30.08/0.8858 35.27/0.9540
SRFormer [82] 10.7 35.02/0.9323 30.94/0.8540 29.48/0.8156 30.04/0.8865 35.26/0.9543
ESC (FT) 12.5 35.14/0.9330 31.10/0.8552 29.53/0.8167 30.23/0.8895 35.60/0.9555
SwinIR [36] ×$\times 4$ 11.9 32.92/0.9044 29.09/0.7950 27.92/0.7489 27.45/0.8254 32.03/0.9260
EDT-B [33] 11.6 33.06/0.9055 29.23/0.7971 27.99/0.7510 27.75/0.8317 32.39/0.9283
CAT-A [9] 16.6 33.08/0.9052 29.18/0.7960 27.99/0.7510 27.89/0.8339 32.39/0.9285
ART [71] 16.6 33.04/0.9051 29.16/0.7958 27.97/0.7510 27.77/0.8321 32.31/0.9283
ACT [67] 46.0 32.97/0.9031 29.18/0.7954 27.95/0.7507 27.74/0.8305 32.20/0.9267
SRFormer [82] 10.6 32.93/0.9041 29.08/0.7953 27.94/0.7502 27.68/0.8311 32.21/0.9271
ESC (FT) 12.5 33.00/0.9054 29.21/0.7968 27.95/0.7504 27.89/0.8351 32.54/0.9295
11 Classic SR Results on Larger Model Size
Although we have demonstrated substantial performance gains over lightweight Transformers (Params$<$1M), larger models (Params$>$10M) remain an active area of research. To assess our method in this regime, we scale ESC to 12.5M parameters, on par with the size of SwinIR, and train and evaluate it accordingly. The scaled ESC uses the window size of 48$\times$48, $N=8$, $M=5$, $C=192$, $C_{\mathrm{ConvAttn}}=48$, and $h=24$. Extra layer normalizations are placed before the $\mathrm{ConvFFN}$s. We leverage the DF2K dataset and follow the ATD’s training strategy [73], pre-training on small patches (64$\times$64) and then fine-tuning on larger patches (96$\times$96). Table 10 shows that ESC delivers performance on par with other large-scale SR Transformers.